# KVCrush: Key Value Cache size-reduction using similarity in head-behaviour

### v1

@@ Abstract
Key-value (KV) caching has emerged as a crucial optimization technique for accelerating inference in large language models (LLMs). By allowing the attention operation to scale linearly rather than quadratically with the total sequence length, KV caching significantly enhances generation throughput [31]. However, due to large context lengths in the modern LLMs, the memory footprint of the KV is a huge bottleneck for model deployment directly impacting the model’s batch size, hindering its ability to deliver high-throughput [18]. Existing research addresses this challenge using several techniques, such as discarding low-attention tokens, quantization, and matrix approximation which typically lead to a negative impact on the model accuracy.
@@
In this paper, We propose KVCrush technology which can be combined with many KV compression technologies to improve the model accuracy at a much smaller memory. KVCrush provides an alternate representation scheme for key-value states, along with a low-overhead token pruning algorithm that accounts for the token distribution in the KV cache, which in turn allows for a a smaller footprint while maintaining the accuracy of the model. Based on our results, KVCrush reduces LongBench [3] KV Cache size by with less than accuracy drop and achieves state-of-the-art average accuracy with minimal overhead, incurring less than total inference latency. KVCrush not only outperforms the accuracy of state-of-the-art importance-based token retention schemes but is also compatible with typical practical LLM deployments using KV cache paging schemes such as vLLM and mixed precision quantization.
I Introduction
@@
Generative AI models, such as large language models (LLMs), have revolutionized the computing industry. These models boast an extensive number of parameters and consistently achieve state-of-the-art performance across various downstream tasks [23] [13] [7] [19]. However, the current trend of model size growth toward multi-trillion parameters [12] — with models growing by one estimate [11] at a staggering rate of every 2 years— poses huge challenges for deploying them in practice. For instance, GPT-175B [4] requires 325GB of memory just to load the model weights. Additionally, not only do inferences for these models strain the platform compute and memory resources (both in terms of bandwidth and capacity), but typically, this is also coupled with strict latency requirements (in the order of tens of milliseconds) which further complicates the problem and poses a significant engineering challenge for efficiently delivering high inference throughput while maintaining low latency for user requests.
@@
LLMs typically consist of stacked transformer decoder layers, with the self-attention module being a critical building block. This module weigh the importance of different tokens capturing their contextual information by attending token pairs at different positions within the input sequence. Nevertheless, self attention matrix multiplication is compute intensive and has a quadratic computational complexity with respect to sequence length[24] which significantly impacts the inference throughput, especially for longer sequences[22].
@@
KV Cache has emerged as the de facto solution to this issue, converting the time complexity of token generation from quadratic to linear, albeit with increased memory overhead proportional to the context length. This mechanism is crucial for autoregressive tasks, where the model generates one output token at a time. As shown in Figure 1, and since in the decode phase each token depends on the key and value tensors of all previous tokens (including the input tokens’ KV tensors computed at prefill, and all new KV tensors computed until the current time step). To avoid recomputing all these tensors for all tokens at each time step, a KV cache is used to cache the values in memory, and for every iteration, when new elements are computed, they are simply added to the running cache to be used in the next iteration.
@@
For instance, consider the OPT-175B model with its impressive 175 billion parameters, which consumes approximately 325 GB of memory. However, when using a batch size of 128 and a sequence length of only 8K, the KV cache requires around 4608 GB of memory. This is an order of magnitude (12X) larger than the model weights themselves. The issue of KV cache size has become increasingly prominent and is a significant cost factor in deploying large language models (LLMs). This is true, especially with the recent trend of LLMs models that have been meticulously developed to scale up to handle extremely long context (for example, Google’s Gemini-pro-1.5 has shown to support a staggering 1M token context length [10]
Several approaches have been proposed to mitigate the memory bottleneck associated with KV caching. Recent research have explored different optimizations of KV caching, including approaches such as low-rank decomposition of the KV cache(e.g., [6]) or pruning non-essential KV cache [31, 16, 18, 9], however, most of these techniques struggle to maintain the accuracy of the model at a smaller KV cache footprint when compared to the full model.
To address this challenge, we introduce KVCrush, a novel KV cache optimization, that provides an alternate representation scheme for key-value states, along with a low-overhead token pruning algorithm that accounts for the token distribution in the KV cache. KVCrush, in turn, allows for a a smaller footprint while maintaining the accuracy of the model and can be easily combined with many KV Cache compression technologies, furthermore, KVCrush also works seamlessly with KV cache paging schemes (such as vLLM [14]) and mixed precision quantization [15] typically used in practical deployments. The technical contributions of this work can be summarized as follows:
@@
Accuracy-aware Low-Overhead KV Optimization Algorithm: Our algorithm leverages the alternative binary representation of tokens and an anchor point to bucketize tokens into representative groups, using very efficient low-overhead distance calculations that scale linearly with the number of tokens. This approach ensures that different token groups are represented in the reduced KV memory footprint and maintaining the inference accuracy of the model.
We show that KVCrush can be efficiently combined with other KV cache optimizations technologies for compression, quantization, and paging making it easily deployable in practice.
We show that for a given memory budget, KVCrush can yield better accuracy than the SOTA KV cache compression schemes alone. We also show KVCrush can achieve better compression than other compression techniques for iso-accurate models, In particular, KVCrush reduces the LongBench [3] KV Cache size by with less than drop in accuracy w.r.t. to the full cache.
We demonstrate that KVCrush is a highly efficient technique, incurring minimal overhead. It achieves state-of-the-art average accuracy on the LongBench, with less than overhead in total inference latency.
II Related Work
@@
II-A Quantizing KV Cache
Quantization is the process of reducing the precision of a model’s parameters and activations to lower bit-widths to save memory and computational resources. Quantization can be categorized into post-training quantization (PTQ) and quantization-aware training (QAT), with PTQ often preferred for large language models due to its lower resource requirements. [26], [17], and [21] demonstrated that quantizing queries, keys, and values to INT8 enables efficient attention operations. However, these methods apply quantization uniformly across all KV cache tokens, and thus will typically negatively impact the model generation accuracy.
On the other hands, research work such as [25], [29], [5], [28] employ mixed precision quantization to allocate different number of bits to various model components or tensors, and thus enable a more efficient compression. This methods leverage the insight that different parts of a model exhibit varying levels of sensitivity to quantization.
As we discuss later in details, KVCrush is running on the KV cache to determine the subset of tokens to retain or to evict. As a result, it is complementary to quantization technology used. The (mixed)-precision used to represent each token can be easily combined with KV-crush by assigning different precision bits to the set of retained tokens based on the "importance" of the token as we will explain.
II-B Sharing KV Cache
Multi-Query Attention (MQA) and Grouped Query Attention (GQA) are techniques developed to address the memory footprint issues LLMs allowing to share KV caches across heads. MQA, introduced by [20], reduces memory usage by sharing key and value representations across all attention heads, which enhances memory efficiency and inference speed, at the expense of generation quality. GQA, proposed by [2], extends this concept by grouping query heads to share KVs, balancing memory reduction with better performance retention. However, GQA involves higher training costs and complexity. Both methods offer trade-offs between memory efficiency and model performance, with MQA favoring memory savings and GQA providing a more balanced approach.
Intuitively, KVCrush is completely orthogonal to these sharing schemes of the KV cache. Our technology is applicable to the deployed KV cache irrespective of whether this cache is shared and grouped across attention heads or not.
II-C Evicting inconsequential Keys and Values
@@
Different strategies for KV pruning and selectively dropping tokens from the KV cahce have been proposed in recent research work, For example, In StreamLLM [27], only the most recent tokens and attention sinks (first few tokens) are retained. H2O [31] and Scissorhands [18] utilize attention-based metrics to determine eviction, with H2O summing attention weights and Scissorhands considering the frequency of attention surpassing a threshold. FastGen [9] combines these methods with heuristics for special tokens. SnapKV [16] similarly depends on attention weights to prune the KV cache, but addresses the efficiency during prefill with long context, and due to the complexity of computing attention score for the whole context, it limits the observation window to the final tokens of the input prompt, thus reducing the complexity from to , where is the context length, while it used max pooling to retain neighboring KVs. PyramidKV [30] builds on SnapKV by configuring variable eviction rates across layers, with more aggressive pruning in later layers where attention is less evenly distributed.
The common drawback across these existing methods is that the model accuracy drops with the higher the KV compression ratio. These works concentrate solely on finding a set of pivotal tokens (the specifics of the algorithms to select these will vary) and evict the rest of the tokens which shrink the KV cache size but at the expense of a negative impact on the model generation quality. KVCrush addresses this drawback by trying to represent the evicted tokens in the final KV cache. It ensures that different token groups are represented in the reduced KV memory footprint, and thus can sustain the accuracy. As a result, KVCrush can be combined with many existing methods to remedy the accuracy drop at the same KV cache budget. We present results in Section IV, combining KVCrush with state-of-the-art research, showing the improved performance of our technology at the same compression ratio.
@@
Experiments | RQ3, RQ5 | RQ1, RQ2, RQ4 | |||
@@
Cache Budget Partitioning | 25% KVCrush + 75% Baseline | Best of (25%+75%) and (75% and 25%) | |||
@@
III KVCrush
The basic flow of KVCrush is shown in Figure 2. Given a certain total budget for the KV Cache, this is split into two smaller portions, represents the cache budget available to store the set of pivotal tokens. This set is determined based on the specifics of KV cache compression algorithm used (e.g. H20, SnapKV, PyramidKV, etc.). While represents the cache budget available to store along a set of representative tokens, these act as proxies of the evicted tokens, and are selected based on low-overhead grouping algorithm (discussed in Section III-B) to ensure better representation of the whole context.
III-A KVCrush Alternative Token Representation
In the KV cache tokens are represented as floating-point vectors of size each. is the size of embedding length and is typically not small. For example, for GPT-3 [4] D is 2048 dimensions while for LLaMA-65B [1] it is set at 4096 dimensions. As previously mentioned, KVCrush will try to group tokens in order to select representative tokens from each group. As a result, it is essential to minimize the overhead of the grouping algorithm. The running time of any clustering algorithm in dimensions will be proportional to the value of .
@@ III-A KVCrush Alternative Token Representation
KVCrush draws some insights of alternative token representation from the FastGen research [9]. The authors in this work demonstrated that distinct attention heads exhibit unique structural patterns and attend to different tokens differently. For example, some attention heads will attend more to special tokens, others will attend more to locality among tokens, others will attend more to punctuation, etc.
Building on these findings we can deduce that the attention score of a certain token across tokens is a vector of length and will represent good semantic information about the contextual properties of that token. Keeping in mind that , where for example, is 96, 128 for GPT-3 [4] and LLaMA-65B [1], respectively.
To reduce the grouping overhead even more, KVCrush takes the alternative token representation one step further. Instead of using a floating point vector of size each, it converts that into a binary vector of size replacing the floating point with a single bit. The main insight here is, for a given head and for any given KV compression algorithm the binary decision of whether to retain or evict the token (i.e. whether this token is pivotal or not) maintains the sematic information of the token importance (a combination of attention score, recency, etc.) with respect to the remaining tokens in the current context.
Figure 3 depicts how KVCrush generates the alternative binary representation of size for the input tokens and it can be summarized in the following steps:
Compute the attention weight matrix using the query and key matrix. The attention score is computed. It should be noted that attention computation has to be done anyway during inference. KVCrush uses the attention values without recomputation.
Apply a threshold per attention head For each attention head the attention weight is normalized and a binary value is used to indicate whether the the token is retained or evicted by the corresponding compression algorithm. For example, the H20 algorithm [31] to achieve compression ratio sets a threshold representing the 50th percentile of attention score on the attention weight matrix. Any tokens with higher attention will be retained (bit = 1) and those with lower attention will be discarded (bit = 0)
Collate one selection bit per head form the binary feature vector for a given token by collating all bits across the different attention heads of the model.
III-B KVCrush Token Grouping and Cache Eviction
As previously mentioned, KVCrush will select tokens that act as proxies of the evicted tokens based to keep in the KV Cache. The selection is based on a low-overhead grouping algorithm built on top of the alternative binary representation of tokens. Figure 4 depicts the main steps of this weak clustering algorithm to form token groups and can be summarized as follows:
@@ III-B KVCrush Token Grouping and Cache Eviction
For each of the input vectors, hamming distance is Computed with respect to the anchor point. The binary representation of tokens makes the hamming distances very low overhead and efficiently computed on the fly using masked comparisons.
Each token is then assigned to one bucket out of set of available. The selected bucket is the one with the the lowest hamming distance. .
After the bucketization of all tokens to their corresponding buckets, One representative vector is selected (in random) from each bucket and all the other () remaining tokens are dropped. Those retained tokens, in addition to the tokens retained by the KV Compression algorithm represent the final tokens of the KV cache.
It should be mentioned however, that we described KVCrush grouping algorithm using Hamming distance computations with respect to ONE anchor point, and thus, token pruning only requires Hamming distance comparisons instead of using distance computations required by standard clustering algorithms. In section IV-C we present results for using a higher overhead clustering algorithm on accuracy and latency of KVCrush.
IV Experiments
We organize our experiments to answer the following questions:
RQ1: What impact do KVCrush settings, such as the cache budget partitioning and anchor points, have on accuracy?
RQ2:How does the accuracy versus latency trade-off of KVCrush compare to traditional clustering algorithms like k-means?
RQ3: What are the accuracy implications of integrating KVCrush with other KV eviction algorithms such as H2O, SnapKV, and Pyramid KV?
RQ4:Is KVCrush effective in paged KV cache settings?
RQ5: How does the accuracy of KVCrush compare to other state-of-the-art KV eviction algorithms?
IV-A Experimental Setup
We summarize experimental setup used for evalution in table I.
IV-A1 Baselines
We compare KVCrush with the following baselines.
FullKV uses the full KV Caches with zero eviction.
H2O [31] uses attention weights sum to determine eviction, keeping top n tokens for each layer.
SnapKV [16] limits the observation window to the final tokens of the input prompt, while using max pooling to retain neighboring KVs.
PyramidKV [30] builds on SnapKV by configuring variable eviction rates across layers, with more aggressive pruning in later layers where attention is less evenly distributed.
IV-A2 Datasets
In our comparative analysis, we benchmarked our results against baselines across 16 datasets from the LongBench suite [3], encompassing a diverse array of tasks and domains such as single-document question answering, multi-document question answering, summarization, few-shot learning, and coding. These datasets feature input lengths ranging from 2K to 18K tokens. Additionally, for other experiments, we utilized the GSM-8K and XSUM datasets from the lm-eval-harness framework [8].
IV-A3 Models
In our evaluation of KVCrush, we employed the LLaMa-3-8B-Instruct and Mistral-7B-Instruct-v0.2 models on the LongBench test suite. Additionally, we used the LLaMa-3-8B-Instruct and Phi-3-mini-4k-instruct models on the lm-eval-harness evaluation framework.

### v2

@@ Abstract
Key-value (KV) caching has emerged as a crucial optimization technique for accelerating inference in large language models (LLMs). By allowing the attention operation to scale linearly rather than quadratically with the total sequence length, KV caching significantly enhances generation throughput (h2o). However, due to large context lengths in the modern LLMs, the memory footprint of the KV is a huge bottleneck for model deployment directly impacting the model’s batch size, hindering its ability to deliver high-throughput (scissorhands). Existing research addresses this challenge using several techniques, such as discarding low-attention tokens, quantization, and matrix approximation which typically lead to a negative impact on the model accuracy.
@@
In this paper, we propose KVCrush technology which can be combined with many KV compression technologies to improve the model accuracy at a much smaller memory. KVCrush provides an alternate representation scheme for key-value states, along with a low-overhead token pruning algorithm that accounts for the token distribution in the KV cache, which in turn allows for a smaller footprint while maintaining the accuracy of the model. Based on our results, KVCrush reduces LongBench longbench KV Cache size by with less than accuracy drop and achieves state-of-the-art average accuracy with minimal overhead, incurring less than total inference latency. KVCrush not only outperforms state-of-the-art importance-based token retention schemes in accuracy, but also integrates seamlessly with quantization, paging, and head-sharing techniques, requiring no retraining or architectural changes.
1 Introduction
@@
Generative AI models, such as large language models (LLMs), have revolutionized the computing industry. These models boast an extensive number of parameters and consistently achieve state-of-the-art performance across various downstream tasks sun2019bert4rec memrec dosovitskiy2020image raffel2020exploring. However, the current trend of model size growth toward multi-trillion parameters isaev2023scaling — with models growing by one estimate gholami2024ai at a staggering rate of every 2 years— poses huge challenges for deploying them in practice. For instance, GPT-175B gpt3 requires 325GB of memory just to load the model weights. Additionally, not only do inferences for these models strain the platform compute and memory resources (both in terms of bandwidth and capacity), but typically, this is also coupled with strict latency requirements (in the order of tens of milliseconds) which further complicates the problem and poses a significant engineering challenge for efficiently delivering high inference throughput while maintaining low latency for user requests.
@@
LLMs typically consist of stacked transformer decoder layers, with the self-attention module being a critical building block. This module weighs the importance of different tokens—capturing their contextual relationships—by attending over key–value token pairs at different positions within the input sequence. Nevertheless, self-attention matrix multiplication is compute intensive and has a quadratic computational complexity with respect to sequence length allyouneed which significantly impacts the inference throughput, especially for longer sequences survey.
@@
KV caching has emerged as the de facto solution to this issue, converting the time complexity of token generation from quadratic to linear, albeit with increased memory overhead proportional to the context length. This mechanism is crucial for autoregressive tasks, where the model generates one output token at a time. In the decode phase, each token depends on the key and value tensors of all previous tokens (including the input tokens’ KV tensors computed at prefill and all new KV tensors computed until the current time step). To avoid recomputing these tensors for every token at each step, a KV cache stores them in memory, incrementally appending new entries to the cache as generation progresses.
@@
For instance, consider the OPT-175B model with its impressive 175 billion parameters, which consumes approximately 325 GB of memory. However, when using a batch size of 128 and a sequence length of only 8K, the KV cache requires around 4608 GB of memory. This is an order of magnitude (12X) larger than the model weights themselves. The issue of KV cache size has become increasingly prominent and is a significant cost factor in deploying large language models (LLMs). This is true, especially with the recent trend of LLM models that have been meticulously developed to scale up to handle extremely long context (for example, Google’s Gemini-pro-1.5 has shown to support a staggering 1M token context length Gemini
Several approaches have been proposed to mitigate the memory bottleneck associated with KV caching. Recent research have explored different optimizations of KV caching, including approaches such as low-rank decomposition of the KV cache(e.g., decomp) or pruning non-essential KV cache h2o; snapkv; scissorhands; fastgen, however, most of these techniques struggle to maintain the accuracy of the model at a smaller KV cache footprint when compared to the full model.
To address this challenge, we introduce KVCrush, a novel KV cache optimization that provides an alternate representation scheme for key-value states, along with a low-overhead token pruning algorithm that accounts for the token distribution in the KV cache. KVCrush, in turn, allows for a smaller footprint while maintaining the accuracy of the model and can be easily combined with many KV cache compression technologies. It is a modular add-on that augments existing methods by preserving contextual diversity, and requires no model retraining or architectural changes. Furthermore, KVCrush also works seamlessly with KV cache paging schemes (such as vLLM vllm) and mixed precision quantization mixed typically used in practical deployments. The technical contributions of this work can be summarized as follows:
@@
Accuracy-aware Low-Overhead KV Optimization Algorithm: Our algorithm leverages the alternative binary representation of tokens and an anchor point to bucketize tokens into representative groups, using very efficient low-overhead distance calculations that scale linearly with the number of tokens. This approach ensures that different token groups are represented in the reduced KV memory footprint, thereby maintaining the inference accuracy of the model.
Practical Efficiency and Integration: We show that KVCrush can be efficiently combined with token eviction-based KV cache compression methods. It achieves state-of-the-art average accuracy on LongBench, with less than inference latency overhead, and offers up to cache reduction with under accuracy drop compared to the full cache.
2 Related Work
@@
2.1 Quantizing KV Cache
Quantization is the process of reducing the precision of a model’s parameters and activations to lower bit-widths to save memory and computational resources. Quantization can be categorized into post-training quantization (PTQ) and quantization-aware training (QAT), with PTQ often preferred for large language models due to its lower resource requirements. xiao2023smoothquant, liu2024llmqat, and sheng2023flexgen demonstrated that quantizing queries, keys, and values to INT8 enables efficient attention operations. However, these methods apply quantization uniformly across all KV cache tokens, and thus will typically negatively impact the model generation accuracy.
On the other hand, research work such as mq1, mq2, mq3, mikv employ mixed precision quantization to allocate different number of bits to various model components or tensors, and thus enable a more efficient compression. This methods leverage the insight that different parts of a model exhibit varying levels of sensitivity to quantization.
As we discuss later in detail, KVCrush operates directly on the KV cache to determine which tokens to retain or evict. Since it does not alter the cache format itself, it is naturally complementary to quantization techniques. In fact, KVCrush can work seamlessly with mixed-precision KV caches by assigning different bitwidths to retained tokens based on their importance.
2.2 Sharing KV Cache
Multi-Query Attention (MQA) and Grouped Query Attention (GQA) are techniques developed to address the memory footprint issues LLMs allowing to share KV caches across heads. MQA, introduced by (mqa), reduces memory usage by sharing key and value representations across all attention heads, which enhances memory efficiency and inference speed, at the expense of generation quality. GQA, proposed by (gqa), extends this concept by grouping query heads to share KVs, balancing memory reduction with better performance retention. However, GQA involves higher training costs and complexity. Both methods offer trade-offs between memory efficiency and model performance, with MQA favoring memory savings and GQA providing a more balanced approach.
Intuitively, KVCrush is completely orthogonal to these KV cache sharing schemes. It operates directly on the deployed cache, regardless of whether key and value states are shared or grouped across attention heads.
2.3 Evicting inconsequential Keys and Values
@@
Different strategies for KV pruning and selectively dropping tokens from the KV cache have been proposed in recent research work, For example, in StreamLLM streamingllm, only the most recent tokens and attention sinks (first few tokens) are retained. H2O h2o and Scissorhands scissorhands utilize attention-based metrics to determine eviction, with H2O summing attention weights and Scissorhands considering the frequency of attention surpassing a threshold. FastGen fastgen combines these methods with heuristics for special tokens. SnapKV snapkv similarly depends on attention weights to prune the KV cache, but addresses the efficiency during prefill with long context, and due to the complexity of computing attention score for the whole context, it limits the observation window to the final tokens of the input prompt, thus reducing the complexity from to , where is the context length, while using max pooling to retain neighboring KVs. PyramidKV pyramidkv builds on SnapKV by configuring variable eviction rates across layers, with more aggressive pruning in later layers where attention is less evenly distributed. DynamicKV dynamickv adaptively resizes KV caches across layers based on task type. LaCache lacache introduces a ladder-shaped caching pattern that prioritizes temporal coverage across layers.
The common drawback across these existing methods is that model accuracy drops as the KV compression ratio increases. These works focus solely on selecting a set of pivotal tokens (using varying algorithms) and evict the rest, which reduces KV cache size but at the cost of generation quality. KVCrush addresses this by explicitly representing groups of evicted tokens via a small set of representative tokens, preserving context diversity and mitigating accuracy loss under high compression. It ensures that different token groups are reflected in the reduced KV memory footprint, thereby sustaining accuracy. As a result, KVCrush can be combined with many existing methods to remedy accuracy drop at the same KV cache budget. We present results in Section 4, combining KVCrush with state-of-the-art techniques and demonstrating improved performance at equal compression ratios.
@@
| Experiments | Section 4.2, Section 4.3.2 | Section 4.3.1, Section 4.3.3, Section 4.3.4, Section 4.3.5 | |||
@@
| Cache Budget Partitioning | 25% KVCrush + 75% Baseline | 25% KVCrush + 75% Baseline | |||
@@
3 KVCrush
The basic flow of KVCrush is shown in Figure 1. Given a certain total budget for the KV Cache, this is split into two smaller portions, represents the cache budget available to store the set of pivotal tokens. This set is determined based on the specifics of KV cache compression algorithm used (e.g. H20, SnapKV, PyramidKV, etc.). While represents the cache budget available to store along a set of representative tokens, these act as proxies of the evicted tokens, and are selected based on low-overhead grouping algorithm (discussed in Section 3.2) to ensure better representation of the whole context.
3.1 KVCrush Alternative Token Representation
In the KV cache tokens are represented as floating-point vectors of size each. is the size of embedding length and is typically not small. For example, for GPT-3 gpt3 D is 2048 dimensions while for LLaMA-65B llama it is set at 4096 dimensions. As previously mentioned, KVCrush will try to group tokens in order to select representative tokens from each group. As a result, it is essential to minimize the overhead of the grouping algorithm. The running time of any clustering algorithm in dimensions will be proportional to the value of .
@@ 3.1 KVCrush Alternative Token Representation
KVCrush draws some insights of alternative token representation from the FastGen research fastgen. The authors in this work demonstrated that distinct attention heads exhibit unique structural patterns and attend to different tokens differently. For example, some attention heads will attend more to special tokens, others will attend more to locality among tokens, others will attend more to punctuation, etc.
Building on these findings we can deduce that the attention score of a certain token across heads is a vector of length and will represent good semantic information about the contextual properties of that token. Keeping in mind that , where for example, is 96, 128 for GPT-3 gpt3 and LLaMA-65B llama, respectively.
To reduce the grouping overhead even more, KVCrush takes the alternative token representation one step further. Instead of using a floating point vector of size each, it converts that into a binary vector of size replacing the floating point with a single bit. The main insight is that, for a given head and attention score based token eviction algorithm, the binary decision of whether to retain or evict a token encodes important contextual signals (e.g., attention score, recency), which we use to build token-level semantic fingerprints.
Algorithm 1 depicts how KVCrush generates a binary representation of size for each input token. The process can be summarized in the following steps:
Compute the attention weight matrix: For each head , compute the attention weight matrix using the query and key projections. Note that this attention computation is already performed during inference, and KVCrush leverages these values without recomputation.
Apply a per-head threshold: For each token and head , compute the row-wise normalized attention score . A token is marked with a bit value of 1 if , and 0 otherwise. The threshold is chosen such that it retains a target fraction of tokens for each head, based on desired compression ratio.
Form the binary representation: For each token , collate the bit values across all heads into a binary vector , which captures its cross-head importance signature.
3.2 KVCrush Token Grouping and Cache Eviction
As previously mentioned, KVCrush will select tokens that act as proxies of the evicted tokens based to keep in the KV Cache. The selection is based on a low-overhead grouping algorithm built on top of the alternative binary representation of tokens. Algorithm 2 depicts the main steps of this weak clustering algorithm to form token groups and can be summarized as follows:
@@ 3.2 KVCrush Token Grouping and Cache Eviction
For each of the input vectors, hamming distance is computed with respect to the anchor point. The binary representation enables computing Hamming distances on-the-fly using lightweight bitwise operations, making the grouping process extremely efficient compared to standard clustering.
Each token is then assigned to one of the buckets. The selected bucket is the one with the lowest hamming distance.
After the bucketization of all tokens into their corresponding buckets, one representative vector nearest to the centroid of each bucket is selected, and the remaining tokens are dropped. The retained tokens, along with the tokens preserved by the KV compression algorithm, form the final tokens of the KV cache.
After assigning all tokens to their respective buckets, one representative vector is selected from each by finding the token whose binary representation is closest to that bucket’s centroid. All the other () tokens are dropped. Those retained tokens, in addition to the tokens retained by the KV compression algorithm, represent the final tokens of the KV cache.
It should be mentioned however, that we described KVCrush grouping algorithm using Hamming distance computations with respect to ONE anchor point, and thus, token pruning only requires Hamming distance comparisons instead of using distance computations required by standard clustering algorithms. In Section 4.3.3 we present results for using a higher overhead clustering algorithm on accuracy and latency of KVCrush.
4 Experiments
We organize our experiments to evaluate KVCrush along four dimensions: accuracy and latency trade-offs under constrained cache budgets; ablation studies on anchor selection, budget partitioning, and clustering methods (including comparisons with k-means); integration with existing KV eviction methods such as H2O, SnapKV, and PyramidKV; and performance in paged KV cache settings.
4.1 Experimental Setup
We summarize experimental setup used for Evaluation in table 1.
4.1.1 Baselines
For baseline comparison, we evaluate KVCrush alongside four methods: FullKV uses the complete KV cache with zero eviction. H2O (h2o) determines eviction based on the cumulative attention weights, retaining the top- tokens per layer. SnapKV (snapkv) restricts the observation window to the final tokens of the prompt while employing max pooling to preserve neighboring KVs. PyramidKV (pyramidkv) extends SnapKV by applying variable eviction rates across layers, pruning more aggressively in later layers where attention distributions are less uniform.
4.1.2 Datasets and Models
We evaluate KVCrush on 16 diverse LongBench datasets and the GSM-8K and XSUM tasks from lm-eval-harness, using LLaMa-3-8B-Instruct, Mistral-7B-Instruct-v0.2, and Phi-3-mini-4k-instruct models. Details of datasets, models, and input lengths are summarized in Table 1.
4.2 End-to-End Evaluation
4.2.1 Accuracy